# Initial data analysis of the national German transplantation registry with a focus on kidney transplantation

### v1

@@ Abstract
Before data is used to answer research questions, it is important to conduct an initial data analysis (IDA). The IDA is used to assess the quality of the data, to identify potential problems, and to determine the data’s ability to answer the research question. The STRATOS initiative provides guidelines for an IDA Baillie2022. In this paper, we focus on their recommendations, especially rules three, six, seven, and ten. The third rule is to ensure that the IDA is made transparent and reproducible. The sixth rule states that the IDA should provide clear visualizations and summaries of the data. In the seventh rule, the authors state that the IDA should analyze missing data. Finally, the tenth rule describes that the IDA results should be considered early Baillie2022.
@@
The global median prevalence of CKD is estimated at 9.5% and increases with the income level of a country Bello2024. In western Europe, it is estimated at 10.63%. Long-term treatment options for ESKD include dialysis and kidney transplantation Bello2024. For treated ESKD, the global median prevalence is estimated at 823 per million population (pmp) and in western Europe at 1,034 pmp Bello2024. Compared to dialysis, kidney transplant recipients show a higher quality of life and a lower mortality rate. These benefits are especially pronounced when a transplantation is performed before starting dialysis Amaral2016.
The availability of kidney transplantation as a treatment, however, is greatly limited by the number of available organs. In Germany in 2023, 2,617 registrations to the waiting list for a kidney transplantation were made DSO2023. In the same year, 2,122 kidney transplantations were performed. However, at the end of 2023, 10,454 patients were still on the waiting list DSO2023. Before a kidney transplantation can be performed, compatibility between the donor and recipient needs to be assured to prevent graft rejection and other complications. To optimize treatment effects both on a individual and a population level, an allocation algorithm is used to find recipients for a donor organ. Data of TxReg allow e.g. important insight into the consequences of such organ allocation rules for certain patient groups Kolbrink2024.
We start with notations and definitions including a description of the data structure of the TxReg data in Section Notations and Definitions. We then describe the methods used for an analysis of missing data, data consistency and distribution of event time data in Section Methods. After presenting the results in Section Results we end with a discussion and conclusion in Section Discussion and Section Conclusions.
@@ Methods
The organization that is managing the TxReg data techspec internally uses four relational databases to collect, store, track and export data. For the export (to which we refer as the TxReg data), the data is provided in the form of a relational database, exported as comma-separated values (CSV) files. In the official documentation of the TxReg, three entities are described: transplantations, recipients and donors. Each of these entities in the database has a unique identifier provided by ET. In the Supplementary Table 1, Additional File 1, we provide an overview of the tables in the TxReg data and to which subject types they refer.
@@
All tables can contain missing values. However, especially for qualitative test results, explicit values like “not tested” or “unknown” exist, which we do not consider as missing values in the context of an IDA.
Three data patterns (namely multi-sourced variables, relating observations and data providers and the mixing of cross-sectional and longitudinal data) for our IDA, we describe it in the next subsections.
@@
Each column $\mathbf{X}_{j}$, $j=1,...,p_{\mathbf{X}}$, of table $\mathbf{X}$ is contributed by one of the data providers: ET, DSO or IQTIG. Supplementary Table 5, Additional File 1 shows to which tables which data provider contributed data. Thereby, not all columns may represent unique information, due to the data collection process through three data providers. For example, the donor’s weight is stored in three columns, one from each data provider. We refer to such variables as “multi-sourced variables”. These columns share the same name as prefix and give the data provider identifier as a suffix. Multi-sourced variables can yield conflicting values but can also provide a chance to impute missing values.
@@
The initial data analysis is focusing on kidney transplantations where 1. the recipient is adult, 2. it is the first transplantation for this recipient and 3. the organ is from a deceased donor. The target population was constructed using the ET-reported recipients, donors, and transplantations in T_Recipient, T_Transplantation, T_Kidney_Waitlist, and T_Donor_PM. Recipients with multiple transplantations were identified and removed using T_Transplantation itself, but also using T_Kidney_Waitlist.C_KidneyTXCount. We only considered columns that had any data on this target population and removed duplicated columns, which led to the final set of 1123 columns used in the analyses. More details on the filtering process can be found in the supplementary material under Supplementary Information 1, Additional File 1. Our population consisted of 14,954 recipients and 9,964 donors across 25 tables.
@@
Usually, the proportion of missing values is computed as the number of missing values divided by the total number of observations $n_{\mathbf{X}}$. In our setting, this denominator is inappropriate, because aspects of the data representation induce structural missingness that does not reflect data quality. In addition to the mixed cross-sectional/longitudinal structure described in Section Cross-Sectional and Longitudinal Data, further representation issues occur: information from different data providers for the same event is usually stored in a single row with data-provider-specific columns, but in some instances it is stored across multiple rows. Consequently, some entries are systematically missing by design (representation/storage) rather than due to incomplete reporting.
@@
In a complex dataset, not only the amount of missing data, but also the structure of the missing data is relevant for analysis. It can reveal mechanisms and potential sources for biases. Two methods are used to analyze the structures in the missing data. First, influx and outflux analysis according to vanBuuren2018 is performed to identify patterns of missing data. Second, to identify potential causal mechanisms for missing data, a decision tree analysis to understand the structure of the missing data is conducted based on the work by Tierney2015. The first approach works vertically on the columns, while the second one operates horizontally on the observations.
@@
The target variable for training the trees is the proportion of missing values per observation. We calculate trees separately by data provider for the reasons described in Section Relating Observations to Data Providers. Thus, for a data provider DP the target variable for observation $i\in\mathcal{I}_{\mathbf{X}}(\text{DP})$ is calculated as:
@@
Beyond data completeness, data consistency represents a critical aspect for reliable analyses. In particular, the multi-sourced variables, as defined in Section Multi-Sourced Variables, present both analytical challenges and opportunities. While they cause collinearity that negatively affects analyses e.g. by increasing standard errors, imputation strategies could highly benefit from using this information: Missing values in one data provider’s column can potentially be imputed using values from another data provider’s corresponding column, thereby improving overall data completeness. We therefore first examine the potential of multi-sourced variables for imputation and second assess the degree of agreement between information reported by several data providers. These analyses can inform both decisions on data preprocessing and variable selection.
The proportion of usable cases $U(\mathbf{X}_{j},\mathbf{X}_{k})$ (see Section Influx and Outflux Analysis) will be used to identify the potential of missing data imputation by the multi-sourced variables. To analyze the consistency and agreement of information between columns, Cramér’s V and the Pearson correlation coefficient are calculated. Cramér’s V is a measure of association between two categorical variables and is calculated from the chi-squared statistic. It ranges from 0 (no association) to 1 (perfect association). For two numerical variables, the absolute value of the Pearson correlation is calculated, which ranges from 0 (no linear association) to 1 (perfect linear association). The absolute value is used as we are interested in the strength of the association.
@@
For event time analyses of graft and patient survival, the following information is needed: The date of organ transplantation, the date of patient death, the date of organ failure and the date of last contact. The latter is required to properly consider the time patients are at risk and observed for experiencing events in estimates of incidence rates and proportions. As already mentioned dates cannot be interepreted as calendar dates (see Section Notations and Definitions), instead dates will refer to relative dates only.
@@ Results
Figure 1 illustrates the proportion of missing data in each column of the tables $PM(\mathbf{X}_{j})$, as defined in Section Missing Data analysis.
@@
These findings show that simple ad-hoc imputation strategies like mean imputing e.g. should not be applied. We recommend the proper separation of the different types of observations (e. g. lab test types) before imputation. The following results on a missing data structure analysis as described in Section Missing Data Structure Analysis can support researchers in defining a reasonable imputation strategy:
Good candidates for an imputation model are column pairs which have non-missing values co-occurring with missing values or vice versa. These two properties are measured with the outflux $O(\mathbf{X}_{j})$ and influx $I(\mathbf{X}_{j})$, as described in Section Influx and Outflux Analysis. The results are shown in Figure 2.
@@
Missing values might be caused by known or unknown mechanisms in the data collection and reporting process. A better understanding of these mechanism can help e.g. to decide whether missing data are missing at random and could inform the construction of imputation models. To identify these mechanisms, in the following results of the missing data structure analysis described in Section Missing Data Structure Analysis. are shown. Table LABEL:tbl-inst-missing-rmse shows columns which were important predictors for the proportion of missing values per observation $OPM(\mathbf{X},\text{DP},i)$. The RMSE gives the prediction error on the test set. Because analyses were performed separately for each DP, results are given for each $\mathbf{X}$ and DP. Columns with an adjusted importance above 50% were considered an important predictor and are shown in this table. Full results are available in Supplementary Table 4, Additional File 1.
@@
Center identifiers were also prominent predictors, appearing in three table-data-provider pairs. In T_Recip_Immuno, the diagnostic center C_DiagCenter was the second most important predictor alongside the test type, indicating that different centers report different sets of results. For T_Transplantation and T_Kidney_Followup, various center columns (follow-up center, registration center, transplant center) predicted missingness, suggesting center-specific reporting practices. Here, the centers might perform different types of follow-ups or have varying protocols for data collection. This should be considered when defining imputation strategies.
@@
First, we analyzed the multi-sourced variables (see Section Multi-Sourced Variables). Figure 3 visualizes the number of columns that represent these multi-sourced variables, grouped by the data provider.
@@
There are multiple sources for the date of last follow-up in the dataset. First, ET directly provides a date for the last follow-up in T_Transplantation.C_LastDate. In the following, we will refer to this date as the “Reported Last Follow-Up Date” (“Reported LFUD”). Furthermore, ET and IQTIG each provide a series of $0,1,\ldots$ follow-up dates in multi-sourced variable columns T_Kidney_Followup.C_Date. The corresponding combined last follow-up date can be defined as the latest date for each recipient. We call this the “Derived LFUD”. Data provided by the DSO includes only donor data without recipient follow-up.
@@
By only considering dates of death and graft failure within the first 3 years after transplantation, we consider the findings of Section Distribution of Event Time Data indicating that information beyond year 3 may be unreliable.
@@
Kaplan-Meier estimates $\hat{S}$ are calculated separately for each of the three outcome definitions. Results are shown in Figure 9.
Figure 9: Kaplan-Meier estimates of the survival probability over time.
@@
  • Conversely, when relying mainly on information provided by IQTIG (“b) IQTIG Reported”), higher survival probabilities are estimated compared to the results using information from ET. This can be explained by the fact that many patient deaths and graft failures are not reported by IQTIG (see Section Time to Patient Death and Section Time to Graft Failure).
@@ Discussion
Based on the finding of our initial data analysis, we propose the following recommendations for data preprocessing and analysis: 1. Tables containing a mix of cross-sectional and longitudinal data should be split by observation type prior to any imputation. 2. Flux analysis can be used to preselect variables for imputation, thereby helping to prioritize and focus efforts. 3. Multi-sourced variables should be used for imputing missing data. For important exceptions as indicated in this paper, we recommend to first harmonize units and codings before imputation using plausible value ranges and mappings. 4. Data from Eurotransplant (ET) and the German Organ Transplantation Foundation (DSO) should be prioritized for primary analyses, while data from the Institute for Quality Assurance and Transparency in Health Care (IQTIG) may serve as a supplementary source for imputation. 5. When defining outcomes for time-to-event analysis, we recommend using a combined outcome that integrates information from all three data providers. Relying solely on data from a single source may lead to biased results.
@@
Other studies using the TxReg data have employed different strategies for extracting the relevant information. vonSamsonHimmelstjerna2023 analyzed patient time on the waiting list for a first kidney transplantation in the “legacy data” between 2006 and 2016. While they reported using data from ET and DSO, they did not mention the use of IQTIG data. Their study population included 24,150 recipients, which is larger compared to our population of 14,954 recipients. This is probably due to different inclusion criteria. Kolbrink2024 combined both “legacy data” and “new data” to analyze the impact of the 65-year age criterion for the “ET Senior Program” (ESP) versus the standard ET Kidney Allocation System (ETKAS) on time spent on the kidney transplant waiting list. They also did not reference the use of IQTIG data. Otto2024 focused on data quality in the “legacy data”, also highlighting challenges related to multi-sourced variables and emphasizing that the different data providers collect data at different points in the transplantation and follow-up process. Their analysis, which covered all organ types, noted that not all recipients, donors, and transplantations are represented consistently across all tables. In our study, we addressed this issue by limiting the target population to recipients with consistently documented transplantations, donors, and follow-ups across all relevant sources (see Supplementary Figures 1 to 3, Additional File 1). Before IQTIG was responsible for collecting the quality assurance data, the BQS Institute had similar role. In studies using DSO data from 2006 and 2008 - linked to BQS data - challenges were reported similar to those we observed in the registry Kutschmann2013; FischerFrhlich2015; Frhauf2011.
@@
Our study has limitations. We focused exclusively on kidney transplant data from the “legacy data” and did not include other organ types or the “new data”. The “new data” differs in structure and includes information collected after 2016, which could offer longer follow-up times and may reflect data patterns following changes in legislation. Additionally, certain data challenges were beyond the scope of this study, such as the need for fuzzy matching of donors and recipients across tables, particularly in the context of repeat transplantations. Kaplan-Meier estimates were used only for illustrative purposes to show the importance of carefully defining event-time outcomes based on TxReg data. A proper analysis of such outcomes should rely on competing risk methods, as recently emphasized by Coemans2025. Although we addressed important issues, we could not give a complete view of the TxReg data in our analysis, due to the complexity of the data.

### v2

@@ Abstract
Before data is used to answer research questions, it is important to conduct an initial data analysis (IDA). The IDA is used to assess the quality of the data, to identify potential problems, and to determine the data’s ability to address the planned research objectives. The STRATOS initiative provides guidelines for an IDA. In this paper, we follow these recommendations, with an emphasis on transparency and reproducibility, clear visualizations and summaries of the data, a systematic assessment of missing data, and early consideration of IDA findings to inform subsequent analyses Baillie2022.
@@
Long-term treatment options for ESKD include dialysis and kidney transplantation Bello2024. Compared to dialysis, kidney transplant recipients show a higher quality of life and a lower mortality rate. These benefits are especially pronounced when a transplantation is performed before starting dialysis Amaral2016. The availability of kidney transplantation as a treatment, however, is greatly limited by the number of available organs. In Germany in 2023, 2,617 registrations to the waiting list for a kidney transplantation were made DSO2023. In the same year, 2,122 kidney transplantations were performed. However, at the end of 2023, 10,454 patients were still on the waiting list DSO2023. Before a kidney transplantation can be performed, compatibility between the donor and recipient needs to be assured to prevent graft rejection and other complications. To optimize treatment effects both on a individual and a population level, an allocation algorithm is used to find recipients for a donor organ. Data from the TxReg allow e.g. important insight into the consequences of such organ allocation rules for certain patient groups Kolbrink2024.
When analyzing outcomes of kidney transplantation, there are different endpoints, which can be derived. When looking at graft failure, the following endpoints are common EBPG2002:
  1. 1.
    Patient survival: Time from transplantation to death; individuals alive at last follow-up are censored at that time.
  2. 2.
    Graft and patient survival: Time from transplantation to irreversible graft failure (return to long-term dialysis or retransplantation) or death.
  3. 3.
    Death-censored graft survival: Time from transplantation to irreversible graft failure, with death while the graft is still functioning treated as a censoring event at death.
We start with notations and definitions including a description of the data structure of the TxReg data. We then describe the methods used for an analysis of missing data, data consistency and distribution of event time data. After presenting the results, we end with a discussion and conclusion.
@@ Methods
The organization that is managing the TxReg data techspec internally uses four relational databases to collect, store, track and export data. For the export (to which we refer as the TxReg data), the data is provided in the form of a relational database, exported as comma-separated values (CSV) files. In the official documentation of the TxReg, three entities are described: transplantations, recipients and donors. Each of these entities in the database has a unique identifier provided by ET. In the Supplementary Table 1, Additional File 1, we provide an overview of the tables in the TxReg data and to which subject types they refer. Linkage between those entities is possible via the pseudonymized identifiers provided by ET txregHandbuch2025.
@@
All tables can contain missing values. However, especially for qualitative test results, explicit values like “not tested” or “unknown” exist, which we do not consider as missing values in the context of an IDA. We acknowledge that, depending on the variable and clinical context, such categories may in some settings be treated as missing values. In this analysis, however, we consistently treat them as informative categories rather than missing data.
Three data patterns (namely multi-sourced variables, relating observations and data providers and the mixing of cross-sectional and longitudinal data) play an important role for our IDA, we describe it them in the next subsections.
@@
Each column $\mathbf{X}_{j}$, $j=1,...,p_{\mathbf{X}}$, of table $\mathbf{X}$ is contributed by one of the data providers: ET, DSO or IQTIG. Supplementary Table 5, Additional File 1 shows to which tables which data provider contributed data. Thereby, not all columns may represent unique variables, due to the data collection process through three data providers. For example, the donor’s weight is stored in three columns, one from each data provider. We refer to such variables as “multi-sourced variables”. Their columns share the same name as a prefix and give the data provider identifier as a suffix. Multi-sourced variables can yield conflicting values but can also provide a chance to impute missing values.
@@
For example, in T_Recipient, $\mathcal{C}_{\mathbf{X}}(\text{ET})$ contains all columns with suffix ET (e.g., EBasisGewichtWertET, EBasisGroesseWertET), while $\mathcal{C}_{\mathbf{X}}(\text{IQTIG})$ contains the corresponding IQTIG-reported columns (e.g., EBasisGewichtWertIQTIG, EBasisGroesseWertIQTIG). A complete list of all multi-sourced variable pairs is provided in the digital supplementary material.
@@
In practical terms, an observation can belong to more than one observation set $\mathcal{I}_{\mathbf{X}}(\text{DP})$ (for different values of DP), depending on which provider-specific columns are non-missing in that observation. For example, in T_Transplantation, an observation with non-missing values in both ET and IQTIG columns would belong to both $\mathcal{I}_{\mathbf{X}}(\text{ET})$ and $\mathcal{I}_{\mathbf{X}}(\text{IQTIG})$. Therefore we consider provider-specific “sub-tables” within each TxReg table: for a given table, we analyze ET, DSO, and IQTIG contributions separately. Columns will always belong to a specific provider through $\mathcal{C}_{\mathbf{X}}(\text{DP})$, while observations can belong to multiple providers through $\mathcal{I}_{\mathbf{X}}(\text{DP})$. This directly reflects the TxReg data structure and avoids counting structurally absent entries as missing values.
@@
We are using the “legacy data” (Altdaten), comprising data collected between 2006 and 2016, i.e., before the TPG amendment 2016, which introduced the requirement for informed consent for data collection for research purposes.
The TxReg data primarily cover donors, transplantations, and recipients in Germany. Within Eurotransplant, however, a small number of postmortem donated organs from Germany are transplanted abroad, and conversely, a small number of organs donated abroad are transplanted in Germany. Geographically, the TxReg data therefore include donors from Germany, transplantations in Germany, and recipients in Germany, with minor cross-border exchanges.txregHandbuch2025
The initial data analysis is focusing on kidney transplantations where 1. the recipient is adult, 2. it is the first transplantation for this recipient and 3. the organ is from a deceased donor. The target population was constructed using the ET-reported recipients, donors, and transplantations in T_Recipient, T_Transplantation, T_Kidney_Waitlist, and T_Donor_PM. Recipients with multiple transplantations were identified and removed using T_Transplantation itself, but also using T_Kidney_Waitlist.C_KidneyTXCount. We only considered columns that had any data on this target population and removed duplicated columns, which led to the final set of 1123 columns used in the analyses. More details on the filtering process can be found in the supplementary material under Supplementary Information 1 and Supplementary Figure 1, Additional File 1. Our population consisted of 14,954 recipients and 9,964 donors across 25 tables.
@@
Usually, the proportion of missing values is computed as the number of missing values divided by the total number of observations $n_{\mathbf{X}}$. In our setting, this denominator is inappropriate, because aspects of the data representation induce structural missingness that does not reflect data quality. In addition to the mixed cross-sectional/longitudinal structure described, further representation issues occur: information from different data providers for the same event is usually stored in a single row with data-provider-specific columns, but in some instances it is stored across multiple rows. Consequently, some entries are systematically missing by design (representation/storage) rather than due to incomplete reporting.
@@
For ET-contributed columns in a table, the denominator is the number of observations with at least one non-missing ET value (not all rows of that table) In a complex dataset, not only the amount of missing data, but also the structure of the missing data is relevant for analysis. It can reveal mechanisms and potential sources for biases. Two methods are used to analyze the structures in the missing data. First, influx and outflux analysis according to vanBuuren2018 is performed to identify patterns of missing data. Second, to identify potential causal mechanisms for missing data, a decision tree analysis to understand the structure of the missing data is conducted based on the work by Tierney2015. The first approach works vertically on the columns, while the second one operates horizontally on the observations.
@@
The target variable for training the trees is the proportion of missing values per observation. We calculate trees separately by data provider for the reasons described. Thus, for a data provider DP the target variable for observation $i\in\mathcal{I}_{\mathbf{X}}(\text{DP})$ is calculated as:
@@
Beyond data completeness, data consistency represents a critical aspect for reliable analyses. In particular, the multi-sourced variables, present both analytical challenges and opportunities. While they cause collinearity that negatively affects analyses e.g. by increasing standard errors, imputation strategies could highly benefit from using this information: Missing values in one data provider’s column can potentially be imputed using values from another data provider’s corresponding column, thereby improving overall data completeness. We therefore first examine the potential of multi-sourced variables for imputation and second assess the degree of agreement between information reported by several data providers. These analyses can inform both decisions on data preprocessing and variable selection.
The proportion of usable cases $U(\mathbf{X}_{j},\mathbf{X}_{k})$ will be used to identify the potential of missing data imputation by the multi-sourced variables. To analyze the consistency and agreement of information between columns, Cramér’s V and the Pearson correlation coefficient are calculated. Cramér’s V is a measure of association between two categorical variables and is calculated from the chi-squared statistic. It ranges from 0 (no association) to 1 (perfect association). For two numerical variables, the absolute value of the Pearson correlation is calculated, which ranges from 0 (no linear association) to 1 (perfect linear association). The absolute value is used as we are interested in the strength of the association.
@@
For event time analyses of graft and patient survival, the following information is needed: The date of organ transplantation, the date of patient death, the date of organ failure and the date of last contact. The latter is required to properly consider the time patients are at risk and observed for experiencing events in estimates of incidence rates and proportions. As already mentioned dates cannot be interpreted as calendar dates, instead dates will refer to relative dates only.
@@ Results
Figure 1 illustrates the proportion of missing data in each column of the tables $PM(\mathbf{X}_{j})$.
@@
These findings show that simple ad-hoc imputation strategies like mean imputing e.g. should not be applied. We recommend the proper separation of the different types of observations (e. g. lab test types) before imputation. The following results on a missing data structure analysis can support researchers in defining a reasonable imputation strategy:
Good candidates for an imputation model are column pairs which have non-missing values co-occurring with missing values or vice versa. These two properties are measured with the outflux $O(\mathbf{X}_{j})$ and influx $I(\mathbf{X}_{j})$. The results are shown in Figure 2.
@@
Missing values might be caused by known or unknown mechanisms in the data collection and reporting process.
Table LABEL:tbl-inst-missing-rmse shows columns which were important predictors for the proportion of missing values per observation $OPM(\mathbf{X},\text{DP},i)$. The RMSE gives the prediction error on the test set. Because analyses were performed separately for each DP, results are given for each $\mathbf{X}$ and DP. Columns with an adjusted importance above 50% were considered an important predictor and are shown in this table. Full results are available in Supplementary Table 4, Additional File 1.
@@
Center identifiers were also prominent predictors, appearing in three table-data-provider pairs. In T_Recip_Immuno, the diagnostic center C_DiagCenter was the second most important predictor alongside the test type, indicating that different centers report different sets of results. For T_Transplantation and T_Kidney_Followup, various center columns (follow-up center, registration center, transplant center) predicted missingness, suggesting center-specific reporting practices. Here, the centers might perform different types of follow-ups or have varying protocols for data collection. However, center identifiers in the exported dataset are anonymized/pseudonymized (in accordance with §15g TPG), so linkage to real transplant centers is not possible under the current data access agreement txregHandbuch2025. This should be considered when interpreting center-level differences in missingness patterns.
@@
First, we analyzed the multi-sourced variables. Figure 3 visualizes the number of columns that represent these multi-sourced variables, grouped by the data provider.
@@
There are multiple sources for the date of last follow-up in the dataset. First, ET directly provides a date for the last follow-up in T_Transplantation.C_LastDate. In the following, we will refer to this date as the “Reported Last Follow-Up Date” (“Reported LFUD”). Furthermore, ET and IQTIG each provide a series of $0,1,\ldots$ follow-up dates in multi-sourced variable columns T_Kidney_Followup.C_Date. The corresponding combined last follow-up date can be defined as the latest date for each recipient. We call this the “Derived LFUD”. In contrast to the single ET source “Reported LFUD”, this derived date is calculated from follow-up records contributed by both ET and IQTIG in T_Kidney_Followup and therefore reflects the latest available tracked follow-up entry in that table. Consequently, “Derived LFUD” can still be earlier than “Reported LFUD” if ET reports a later follow-up date in T_Transplantation.C_LastDate that is not present in follow-up observations. Data provided by the DSO includes only donor data without recipient follow-up. recipient follow-up.
@@
By only considering dates of death and graft failure within the first 3 years after transplantation, we consider the findings indicating that information beyond year 3 may be unreliable.
@@
Kaplan-Meier estimates $\hat{S}$ for the composite endpoint of graft and patient survival are calculated separately for each of the three outcome definitions. Results are shown in Figure 9.
Figure 9: Kaplan-Meier estimates of graft and patient survival probability over time (composite endpoint).
@@
  • Conversely, when relying mainly on information provided by IQTIG (“b) IQTIG Reported”), higher survival probabilities are estimated compared to the results using information from ET. This can be explained by the fact that many patient deaths and graft failures are not reported by IQTIG.
@@ Discussion
Based on the finding of our initial data analysis, we propose the following recommendations for data preprocessing and analysis:
  1. 1.
    Tables containing a mix of cross-sectional and longitudinal data should be split by observation type prior to any imputation.
  2. 2.
    Flux analysis can be used to preselect variables for imputation, thereby helping to prioritize and focus efforts.
  3. 3.
    Multi-sourced variables should be used for imputing missing data. For important exceptions as indicated in this paper, we recommend to first harmonize units and codings before imputation using plausible value ranges and mappings.
  4. 4.
    Data from Eurotransplant (ET) and the German Organ Transplantation Foundation (DSO) should be prioritized for primary analyses, while data from the Institute for Quality Assurance and Transparency in Health Care (IQTIG) may serve as a supplementary source for imputation. This is in line with the role of IQTIG as a quality assurance institution.
  5. 5.
    When defining outcomes for time-to-event analysis, we recommend using a combined outcome that integrates information from all three data providers. Relying solely on data from a single source may lead to biased results.
@@
Other studies using the TxReg data have employed different strategies for extracting the relevant information. vonSamsonHimmelstjerna2023 analyzed patient time on the waiting list for a first kidney transplantation in the “legacy data” between 2006 and 2016. While they reported using data from ET and DSO, they did not mention the use of IQTIG data. Their study population included 24,150 recipients, which is larger compared to our population of 14,954 recipients. This is probably due to different inclusion criteria. Kolbrink2024 combined both “legacy data” and “new data” to analyze the impact of the 65-year age criterion for the “ET Senior Program” (ESP) versus the standard ET Kidney Allocation System (ETKAS) on time spent on the kidney transplant waiting list. They also did not reference the use of IQTIG data. Otto2024 focused on data quality in the “legacy data”, also highlighting challenges related to multi-sourced variables and emphasizing that the different data providers collect data at different points in the transplantation and follow-up process. Their analysis, which covered all organ types, noted that not all recipients, donors, and transplantations are represented consistently across all tables. For our study, detailed methodology documentation for each source database (ET, DSO, IQTIG) was only partially available. We therefore relied on the TxReg handbook and publicly available registry documentation txregHandbuch2025. The limited transparency of source-specific documentation is itself an important limitation of the registry and is consistent with the challenges reported by Otto2024.
In our study, we addressed this issue by limiting the target population to recipients with consistently documented transplantations, donors, and follow-ups across all relevant sources (see Supplementary Figures 2 to 4, Additional File 1). Before IQTIG was responsible for collecting the quality assurance data, the BQS Institute had similar role. In studies using DSO data from 2006 and 2008 - linked to BQS data - challenges were reported similar to those we observed in the registry Kutschmann2013; FischerFrhlich2015; Frhauf2011.
@@
Mortality ascertainment data in the TxReg is based on routine documentation at transplant centers (and/or aftercare providers) and subsequent reporting to the respective data providers (ET and IQTIG) txregHandbuch2025. This is different from for example the SRTR database in the US, which is supplemented by the Social Security Death Master File (SSDMF) srtrDatabase. The usage of relative dates in the TxReg instead of absolute dates, limits the potential analysis to external events like the COVID-19 pandemic. this also prevents robust calendar-time stratification into transplantation eras (e.g., 5-year cohorts). As a result, together with the short follow-up, analyses cannot be directly aligned to contemporary practice periods, which limits immediate clinical interpretability.
For clinicians and patients, the central question is what the TxReg can and cannot reliably support in practice.
Strengths of the TxReg data include: (1) broad national coverage across key steps of kidney transplantation care, (2) complementary information from ET, DSO, and IQTIG that can improve endpoint ascertainment when sources are combined, and (3) the possibility to reconstruct recipient-, donor-, and transplantation-level trajectories when records are consistently linked.
Important limitations include: (1) insufficient documentation of data collection processes and individual variable definitions across providers, which hampers transparent interpretation, (2) inconsistent representation of recipients, donors, and transplantations across and within tables, requiring strict inclusion criteria to ensure comparable data, and (3) missing reliable long-term follow-up data.
Accordingly, this IDA supports three robust practical conclusions: first, TxReg data can be used for clinically relevant kidney-transplant analyses when cohorts are restricted to consistently linked recipient, donor, and transplantation records; second, endpoint ascertainment is more complete when information from ET, DSO, and IQTIG is combined rather than relying on a single source; third, reproducible inference requires explicit preprocessing decisions, including source prioritization, harmonized coding/units, and transparent handling of missingness before model-based analyses.
Beyond these registry-level constraints, our study has additional limitations specific to its scope: we focused on kidney transplantation in the “legacy data” and did not include other organs or the structurally different “new data” (post-2016). Some challenges, including approximate record linkage with repeat transplantations, were outside the present scope. Kaplan-Meier analyses were used illustratively; for formal time-to-event inference in this setting, competing-risk methods are more appropriate, as emphasized by Coemans2025.